OpenStudy (anonymous):

A triangle has one side that measures 4, another side that measures x, and a third side that measures 2 more than x. The perimeter is 12. Which equation would we use to find the value of x? x+x= 12 x+x+ 4 = 12 x+x+ 2 = 12 x+x+ 2 + 4 = 12

OpenStudy (anonymous):

perimeter of triangle= side +side+side simply add all three sides (third side is x+2) and solve a bit you will see

OpenStudy (anonymous):

is it c

OpenStudy (ganpat):

no, perimeter = sum of all sides... in given triangle 1side = 4 2nd side = x 3rd side = x + 2 so their sum = 4 + x + x + 2 So option 4... x + x+ 2 + 4 = 12